OpenStudy (anonymous):

Find f'(x) if f(x)= ∫ 0 to x^(2) (sqrt(1+r^(3)))

hartnn (hartnn):

General formula derivative of integration\( \Large \dfrac{d}{dx}\int \limits_{f(x)}^{g(x)}h(x)dx = h(g(x))g'(x)-h(f(x))f'(x)\)

hartnn (hartnn):

in your case, g(x) = x^2 f(x) = 0 h(x) = \(\sqrt{1+r^3}\)

hartnn (hartnn):

f(x) = 0, so just \(\Large h(g(x))g'(x)\)

hartnn (hartnn):

try now :)

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

welcome ^_^